### v1

# Learnable Total Variation with Lambda Mapping for Low-Dose CT Denoising

Abstract
Although Total Variation (TV) performs well in noise reduction and edge preservation on images, its dependence on the lambda parameter limits its efficiency and makes it difficult to use effectively. In this study, we present a Learnable Total Variation (LTV) framework that couples an unrolled TV solver with a data-driven Lambda Mapping Network (LambdaNet) predicting a per-pixel regularization map. The pipeline is trained end-to-end so that reconstruction and regularization are optimized jointly, yielding spatially adaptive smoothing: strong in homogeneous regions, relaxed near anatomical boundaries. Experiments on the DeepLesion dataset, using a realistic noise model adapted from the LoDoPaB-CT methodology, show consistent gains over classical TV and FBP+U-Net: +2.9 dB PSNR and +6% SSIM on average. LTV provides an interpretable alternative to black-box CNNs and a basis for 3D and data-consistency–driven reconstruction.
Index Terms—  Low-Dose CT, Total Variation, Unrolled Optimization, Lambda Mapping, Denoising
1 Introduction
Reducing radiation exposure without compromising diagnostic accuracy remains a key challenge in low-dose computed tomography (LDCT). Lower X-ray doses introduce strong quantum noise and streak artifacts, degrading both visual and quantitative image quality. Traditional denoising methods such as Total Variation (TV) regularization [rudin1992tv] effectively suppress noise but often oversmooth textures and lose fine anatomical details. Deep convolutional neural networks (CNNs) [zhang2017beyond, jin2017fbpconvnet] can produce sharper results, yet may generate artificial structures and lack physical interpretability.
We propose a Learnable Total Variation (LTV) framework that unifies model-based optimization and deep learning for LDCT denoising. The method embeds a TV prior into an unrolled primal–dual solver [chambolle2011firstorder] and uses a U-Net-inspired LambdaNet to predict a spatially varying regularization map ($\lambda$-map). This coupling enables locally adaptive denoising guided by both iterative TV dynamics and learned spatial priors. The unrolled solver remains fully differentiable, while the learned $\lambda$-map modulates the regularization strength to preserve edges and suppress homogeneous noise. Experiments show that the proposed approach improves PSNR by about 2.9 dB and SSIM by 6% over classical TV and FBP+U-Net [jin2017fbpconvnet], delivering more faithful and interpretable reconstructions at reduced dose.
2 Related Work
Classical TV-based Reconstruction. Total Variation (TV) regularization has been a fundamental tool for inverse problems in medical imaging [rudin1992tv]. It penalizes image gradients to suppress noise while preserving edges. Adaptive variants [liu2012adaptive] improve local flexibility but still rely on hand-tuned parameters and lack robustness across dose levels or anatomies.
Deep Learning Approaches. Data-driven models such as RED-CNN [chen2017redcnn] and FBPConvNet [jin2017fbpconvnet] learn mappings between low- and normal-dose CT images. Although they yield sharp results, the absence of explicit physical constraints may cause hallucinated textures. Hybrid unrolled methods [adler2018learned, aggarwal2019modl] mitigate this by embedding learnable modules within iterative solvers.
Learnable Regularization. Recent works explore spatially varying priors [kofler2023ltv, morotti2025edgeaware] to balance interpretability and flexibility. Our framework extends this line by coupling a differentiable unrolled TV solver with a deep $\lambda$-prediction network, enabling region-adaptive and physically consistent regularization.
Fig. 1: Overview of the proposed Learnable Total Variation (LTV) framework. LambdaNet predicts a spatially adaptive $\lambda$-map that guides an unrolled TV solver (Chambolle–Pock algorithm, $T$ steps) to reconstruct a denoised image. The pipeline is trained end-to-end with a composite loss combining image fidelity and $\lambda$-regularization.
3 Methodology
3.1 Overview
The proposed LTV framework integrates a spatially adaptive regularization network with an unrolled optimization module. As shown in Fig. 1, it consists of (i) LambdaNet, predicting a per-pixel $\lambda$-map, and (ii) an unrolled TV denoiser with $T$ primal–dual updates. Given noisy input $y$, ground truth $x^{\star}$, and reconstruction $\hat{x}$, the pipeline is trained end-to-end with gradients propagated through all solver iterations.
3.2 LambdaNet and Adaptive $\lambda$-Mapping
LambdaNet is a U-Net-style encoder–decoder (four downsampling stages, 64 base channels) with attention-gated skip connections. It employs a multi-scale fusion head, combining logits from three decoder stages (full, 1/2, and 1/4 resolution) via a weighted average. This fused logit map $z$ is passed through a final $\tanh$ activation to produce outputs in [1,1][-1,1]. We rescale $z$ into a controlled, spatially-adaptive regularization strength $\lambda$ using a cosine-ramped upper bound $\lambda_{\max}(t)$:
$s$ $\;=\;\mathrm{clip}\!\left(\tfrac{z+1}{2},\,\epsilon,\,1-\epsilon\right),$
$\lambda$ $\;=\;\lambda_{\min}+s\!\left[\lambda_{\max}(t)-\lambda_{\min}\right],$ (1)
where $\lambda\in[\lambda_{\min},\,\lambda_{\max}(t)]$. The ramp on $\lambda_{\max}(t)$ (e.g., $0.40\!\rightarrow\!1.50$ during the first 15 epochs) prevents premature over-smoothing in early training while enabling a wide dynamic range for spatial selectivity in later epochs.
3.3 Unrolled Primal–Dual TV Denoiser
Given the per-pixel map $\lambda$, we use a Chambolle–Pock-style solver unrolled for T=20T{=}20 iterations. The solver features learnable step sizes $(\tau,\sigma_{d})$ and a relaxation parameter $\theta$, all of which are constrained via softplus and clamping to ensure stable ranges. Let $x^{(k)}$ be the primal image and $p^{(k)}$ the dual variable (gradient field).
Dual update. Dual ascent with a pixel-wise projection onto an $\ell_{2}$ ball defined by $\lambda$:
$p^{(k+1)}\;=\;\Pi_{\|\cdot\|\leq\lambda}\!\Big(p^{(k)}+\sigma_{d}\nabla\bar{x}^{(k)}\Big).$ (2)
Here, $\bar{x}^{(k)}$ is the extrapolated primal variable from the previous step. This projection locally bounds the TV penalty based on LambdaNet’s prediction.
Primal update. TV-driven smoothing is balanced with data fidelity:
$x^{(k+1)}\;=\;\frac{x^{(k)}+\tau\nabla^{\!\top}p^{(k+1)}+\tau\,w_{\text{data}}\,y}{1+\tau\,w_{\text{data}}},$ (3)
where $w_{\text{data}}=1/\sigma_{\text{data}}$ is the data fidelity weight and $\nabla^{\!\top}$ denotes the divergence operator (negative transpose of the gradient $\nabla$).
Relaxation. The iterate is then extrapolated to accelerate convergence:
$\bar{x}^{(k+1)}\;=\;x^{(k+1)}+\theta\big(x^{(k+1)}-x^{(k)}\big),$ (4)
Here, $\theta$ is a learnable over-relaxation parameter that provides a momentum effect, using the update vector ($x^{(k+1)}-x^{(k)}$) to accelerate convergence. This new iterate $\bar{x}^{(k+1)}$ serves as input to the subsequent dual update. Equations (2)–(4) are explicitly unrolled, making the solver fully differentiable.
Ground Truth Noisy TV FBP+U-Net LTV (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fig. 2: Representative LDCT slices comparing noisy, TV, U-Net, and LTV reconstructions.
3.4 Comprehensive Training Objective
Our comprehensive objective, $\mathcal{L}_{\text{total}}$, balances image fidelity with sophisticated priors on the $\lambda$-map as a weighted sum of four synergistic components. The effect of different weightings of these four important subcomponents on performance is beyond the scope of this study and has therefore been deferred to future work.
[…]
3.4.1 Image Fidelity
This component anchors the reconstruction $\hat{x}$ to the ground truth $x^{\star}$:
  • $\mathcal{L}_{\text{MSE}}$: The standard $L_{2}$ loss penalizing pixel-wise intensity errors.
  • $\mathcal{L}_{\text{SSIM}}$: Captures structural similarity and contrast (as $1-\mathrm{SSIM}(\hat{x},x^{\star})$), which $L_{2}$ loss overlooks.
  • $\mathcal{L}_{\text{perc}}$: A VGG16-based Perceptual loss that preserves fine textures and mitigates $L_{2}$-induced over-smoothing.
3.4.2 $\lambda$-Map Spatial Smoothness
To ensure the $\lambda$-map is spatially coherent and artifact-free, we apply two priors:
  • $\mathrm{TV}(\lambda)$: A Total Variation penalty to encourage a piecewise-constant map, preventing erratic oscillations.
  • $\mathcal{L}_{\text{spatial}}$: Complements $\mathrm{TV}(\lambda)$ by directly penalizing $L_{1}$ differences between adjacent $\lambda$ values, enforcing localized smoothing.
3.4.3 Structure-Lambda Alignment
[…]
  • $\mathcal{L}_{\text{align}}$ & $\mathcal{L}_{\text{edge}}$ (Guidance Losses): $\mathcal{L}_{\text{proj}}$ provides sparse gradients (only at the band’s boundaries). These auxiliary $L_{1}$ losses provide a dense signal to stabilize training, pulling $\lambda$ towards scaled ($\mathcal{L}_{\text{align}}$) or normalized ($\mathcal{L}_{\text{edge}}$) versions of the gradient map.
3.4.4 $\lambda$-Map Distributional Regularization
To prevent convergence to a degenerate (e.g., spatially constant) solution, we regulate the map’s overall distribution:
  • $\mathcal{L}_{\text{var}}$ (Variance Loss): Formulated as $\mathcal{L}_{\text{var}}=-\mathrm{std}(\lambda)$, this loss encourages diversity by maximizing the map’s standard deviation, forcing a wide dynamic range.
  • $\mathcal{L}_{\text{ent}}$ (Entropy Loss): This high-weight loss ($w_{\text{ent}}=0.05$) prevents distributional collapse, promoting a complex, information-rich map and ensuring a non-trivial, adaptive strategy.
4 Experiments and Results
Datasets and preprocessing. We created a paired dataset from DeepLesion [deeplesion] slices by simulating 10% dose noise based on the LoDoPaB-CT methodology [lodopab]. About $1000$ axial slices were split $70/15/15$ by patient. All data were normalized to $[0,1]$ and trained on $512{\times}512$ patches with mild augmentations (rotations, flips, intensity shifts). Low-variance patches were resampled to ensure structural content.
Model and training. LambdaNet is a 4-level U-Net with attention gates and multi-scale $\lambda$ heads. The $\lambda$-map is constrained to $\lambda_{\min},\lambda_{\max}(t)]$, where $\lambda_{\max}$ ramps from $0.4$ to $1.5$ over the first 15 epochs. Reconstruction uses a 20-step primal–dual TV solver with learnable step-size parameters $(\tau,\sigma_{d},\theta)$. Training is end-to-end via Adam (LambdaNet: $2{\times}10^{-4}$; solver: $1{\times}10^{-5}$), batch size $8$, gradient clipping, and a ReduceLROnPlateau scheduler. [^1] Our implementation is available at: https://github.com/itu-biai/deep_tv_for_ldct
4.1 Quantitative and Qualitative Analysis
Table 1 shows the PSNR and SSIM results of TV, FBP + U-Net and our proposed LTV method. Raw LDCT inputs average $23.04{\pm}2.13$ dB and $0.704{\pm}0.075$. LTV achieves the highest PSNR and SSIM. It yields a substantial +2.9{+}2.9 dB gain over the FBP+U-Net baseline [jin2017fbpconvnet] and surpasses Classical TV in both metrics, demonstrating a superior noise-to-detail balance.
[…]
Method PSNR (dB) SSIM
Noisy Input $23.04\pm 2.13$ $0.704\pm 0.075$
Classical TV $27.99\pm 1.38$ $0.816\pm 0.124$
FBP+U-Net [jin2017fbpconvnet] $26.48\pm 1.39$ $0.784\pm 0.091$
LTV (ours) $\mathbf{29.36\pm 1.69}$ $\mathbf{0.836\pm 0.123}$
The reconstruction results are shown in Fig. 2. A zoomed-in region was extracted for a closer look with both PSNR and SSIM values attached. LTV preserves fine structures, such as vessel continuity, and reduces streak artifacts—critical for diagnostics—while avoiding the over-smoothing of TV or the anatomical inconsistencies of FBP+U-Net.
Absolute error maps represented in (Fig. 3) show how efficient the denoising performs. FBP+U-Net shows localized high-error zones around complex edges. In contrast, LTV yields lower and more spatially consistent (uniform) error distributions, confirming its adaptive weighting balances detail preservation and smoothing more effectively.
Fig. 3: Absolute error maps. LTV yields smaller and more uniform residuals, whereas FBP+U-Net shows high-error zones near edges and lesions.
4.2 Lambda Adaptivity and Ablation
The learned $\lambda$-maps show a non-uniform, sparse distribution (mean $\approx 0.18$, median $\sim 10^{-3}$). A mild positive correlation with gradient magnitude ($r{\approx}0.15$) confirms the adaptive strategy: LTV increases regularization in flat areas and relaxes it near edges. An ablation study underscores the importance of the $\lambda$-map priors: removing the $\lambda$-TV or entropy regularizers degrades performance by ${\sim}0.6$ dB PSNR and ${\sim}0.01$ SSIM, confirming their stabilizing role in controlling $\lambda$.
5 Conclusion
We proposed a Learnable Total Variation (LTV) framework that integrates a data-driven $\lambda$-mapping network with an unrolled TV solver for low-dose CT denoising. The method adaptively balances noise suppression and structure preservation, outperforming classical TV and U-Net baselines on our test set. Future work will extend the framework toward 3D and physics-informed CT reconstruction.
6 Compliance with Ethical Standards
This study used the publicly available, de-identified DeepLesion [deeplesion] dataset. Our low-dose noise simulation was based on the public methodology of the LoDoPaB-CT challenge [lodopab]. No new data collection was performed; experiments were conducted under the respective dataset licenses, and no additional IRB approval was required for analyses of anonymized public data.
7 Acknowledgments
No external funding was received for this study. The authors declare no relevant financial or non-financial conflicts of interest.